Find the area of the region inside both the rose r =sin(2Φ) and the circle r=cosΦ.The appropriate graph and intervals to the best of your interest to solve this question 
1
Expert's answer
2020-07-06T19:02:43-0400
"\sin(2\phi)=\cos(\phi)"

"2\sin(\phi)\cos(\phi)=\cos(\phi)"

"\cos(\phi)=0\text{ or } \sin(\phi)={1\over 2}"




"A_1={1\over2}\displaystyle\int_{0}^{\pi/2}(\sin(2\phi))^2d\phi="

"={1\over4}\displaystyle\int_{0}^{\pi/2}((1-\cos(4\phi))d\phi="

"={1\over4}\big[\phi-{1\over 4}\sin(4\phi)\big]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}={\pi\over 8}(units^2)"


"A_2={1\over2}\displaystyle\int_{\pi/6}^{\pi/2}\bigg((\sin(2\phi))^2-(\cos\phi)^2\bigg)d\phi="

"={1\over4}\displaystyle\int_{\pi/6}^{\pi/2}\bigg(1-\cos(4\phi)-1-\cos(2\phi)\bigg)d\phi="

"=-{1\over4}\bigg[{1\over 4}\sin(4\phi)+{1\over 2}\sin(2\phi)\bigg]\begin{matrix}
 \pi/2 \\
 \pi/6
\end{matrix}="

"=-{1\over4}(0-{\sqrt{3}\over 8}+0-{\sqrt{3}\over 4})={3\sqrt{3}\over 32}(units^2)"


"Area=2(A_1-A_2)=2({\pi\over 8}-{3\sqrt{3}\over 32})={4\pi-3\sqrt{3}\over 16} (units^2)"

"Area=\dfrac{4\pi-3\sqrt{3}}{16} \text{ square units}"